• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Huber, G.

    Log entry time 16:00:03 on November18,2000

    Entry number 52325

    keyword=Day shift summary

    08:00 - Day shift begins with good 50uA running conditions.

    09:06 - We receive permission to go to 75uA while Hall C is still down. CODA rate jumps from 380 to 540 Hz, with 6% deadtime.

    09:20 - CODA crash. All four ROCs were initially in la-la-land, but eventually recovered without reboot. Did kcoda and new runcontrol.

    11:49 - MCC complains about an excessively high RF trip rate, and requests to reduce our beam to 60uA.

    12:08 - FPP chamber 4 (chan 8.1.16) was tripped off. Some data at the end of run 1400 may be bad. Warning script was disabled at when we changed to 60uA, and forgot to restart, so we had no alarm. So we know that the chambers were still on at 11:50, and at most 15 minutes of data were lost. Sorry.

    12:45 - Due to operator error, all master 8.0 channels tripped off while resetting voltage. Run 1401, which was running during this period, should be ignored. Finally start running again.

    12:56 - It appears that the operators have seen that the machine is stable at 60uA, and so have increased us to 65uA. I'm glad they're paying attention.

    14:47 - Q3 quench immediately after the start of run 1405. Ignore this run. We will have to enter the hall to reset the power supply. Note that Hall B had a magnet quench earlier this afternoon due to ESR problems. None of the halls are up right now, all ultimately related to ESR.

    15:25 - Entered hall to reset Q3 supply. Cycled power supply to 1600 A before setting nominal value. While talking to cryo about the problems, we lost Q1 due to low helium level.

    15:50 - Ed Folts is on site, and enters hall to check everything out. Q1 had a hard quench, and required the key on the PS to be turned off and on to reset the interlock.
    Shift ends with beam re-established to the hall.

    Run Start Stop Events Comments
    1396 07:53 08:47 1.0M
    1397 08:48 09:24 0.83M Run end at CODA crash.
    1398 09:28 10:12 1.0M
    1399 10:13 10:58 1.0M
    1400 10:58 12:08 1.0M FPP ch4 tripped off near end of run.
    1401 Garbage run.
    1402 12:45 13:25 1.0M
    1403 13:26 14:08 1.0M
    1404 14:09 14:46 1.0M
    1405 Garbage run.
    1406 15:58